# Data Authenticity, Consent, & Provenance for AI are all broken: what will it take to fix them?

### v1

@@ Abstract
New capabilities in foundation models are owed in large part to massive, widely-sourced, and under-documented training data collections. Existing practices in data collection have led to challenges in documenting data transparency, tracing authenticity, verifying consent, privacy, representation, bias, copyright infringement, and the overall development of ethical and trustworthy foundation models. In response, regulation is emphasizing the need for training data transparency to understand foundation models’ limitations. Based on a large-scale analysis of the foundation model training data landscape and existing solutions, we identify the missing infrastructure to facilitate responsible foundation model development practices. We examine the current shortcomings of common tools for tracing data authenticity, consent, and documentation, and outline how policymakers, developers, and data creators can facilitate responsible foundation model development by adopting universal data provenance standards.
@@ 1 The Need for Data Provenance
In the last decade, data from across the web, such as news, social media, and encyclopedias, has become a vital resource in general-purpose, generative AI [^1] This work is scoped specifically to large foundation models that are (a) generative, meaning they can create potentially inauthentic text, images or other content, and (b) are typically general-purpose, and therefore tend to scrape vast, diverse distribution of web data, that are under-documented, curated, and checked for consenting use. consumer technologies like GPT-4, Midjourney, and Whisper. These technologies, some with over 100 million weekly users (Malik, 2023), have already begun to catalyze innovation (Brynjolfsson et al., 2018) and scientific inquiry while starting to affect wide swaths of the economy and many everyday consumers (Bommasani et al., 2021). These models are trained on diverse compilations of text, image, and audio data scraped from the web, synthetically generated, or hand-curated. The resulting race to scrape, secure, and mass-produce massive collections of loosely structured data has come with consequences.
Current practices include widely sourcing and bundling data without tracking or vetting their original sources (Longpre et al., 2023a), creator intentions (Chayka, 2023), copyright and licensing status (Bandy & Vincent, 2021), or even basic composition and properties (Dodge et al., 2021). The lack of transparency on and public infrastructure available to track these metadata has led developers into ethical and legal challenges.
@@
These issues have motivated new data infrastructure and frameworks (which we survey in Section 5), to overcome challenges in sourcing training data responsibly. We isolate several tools in the existing ecosystem for foundation model data management, that we identify as commonly used or cited in the foundation model literature. However, while these tools offer promising solutions, we find they neglect key aspects of the problems, lack interoperability with parallel standards/tools, or haven’t yet succeeded in reaching widespread adoption (Section 5.5 and Section 6). A unified framework dedicated to the structured documentation of data properties is needed, which requires action from multiple stakeholders:
@@
To respond to this urgent need, we outline the importance of data provenance for different stakeholders (Section 2), recognition of the problem (Section 3), legal and regulatory considerations (Section 4), and how the current set of solutions to trace data authenticity, consent, and provenance have limitations and trade-offs (Section 5). Finally, we outline recommendations to facilitate a standardized library for tracing critical data features, and address challenges facing informed and responsible AI (Section 6 and Section 7).
This work investigates the shortcomings in existing data infrastructure (Section 5), emphasizes the need for interoperable data documentation standards, and identifies the important properties for the design of new data infrastructure tooling (Section 6 and Table 2).
@@ 2.1 Creators: Protecting Rights and Avoiding Harms
The unsettled legal status of AI training data (Congressional Research Service, 2023) has led to some initial compensation proposals (Cheng, 2023), over which creators have little control. They are also rarely credited as having contributed training data, despite the occasional artist signature or watermark that slips through into generated content (Vincent, 2023a). This lack of data transparency and legal clarity has led to several lawsuits against leading AI companies (Vincent, 2022; Small, 2023; Brittain, 2023; Vincent, 2023b) and to calls from creators and publishers for strong transparency requirements around AI data (Cooke, 2023; David, 2023b; Bommasani et al., 2023b).
@@ 2.2 Developers: Informed Data Use
AI developers have an acute interest in data and its provenance for model performance, model behavior, and anticipate limitations and risks. While model performance generally improves with more data, the quality and diversity of data are also critical factors for reliable performance (Kaplan et al., 2020; Hoffmann et al., 2022). Model behavior tends to emulate the structure and composition of data at both the pretraining and finetuning stages  (Chung et al., 2022; Longpre et al., 2023b). For these reasons, developers curate specialized pretraining corpora for scientific writing (Lewkowycz et al., 2022), code (Li et al., 2022), biomedical content (Singhal et al., 2023) and legal works (Henderson et al., 2022). As such, information about data sources and their properties informs AI model training.
Data documentation, provenance and analysis tooling have proven particularly important for helping practitioners understand very large datasets (Bender & Friedman, 2018; Gebru et al., 2021). Existing datasets suffer from discoverability of issues related to biased or sexual content (David, 2023a)), private data (Nasr et al., 2023), copyright infringing data (Min et al., 2023), or non-commercially restricted data (Heath, 2023). Lastly, reproducibility and scientific progress more broadly are accelerated by data transparency and structured documentation (Kapoor & Narayanan, 2022).
@@ 2.3 Broader Impacts: Reducing Risk and Fighting Bias
Societal risks from AI run the gamut from privacy violations (Maiberg, 2023) and exposure of personally identifying information, to systemic economic impacts and job displacement (Autor et al., 2022; Gruetzemacher et al., 2020), to bias and discriminatory behavior (Kurita et al., 2019). These risks are fundamentally tied to the data that models are trained on (Longpre et al., 2023b), which, together with the context and affordances of their applications, dictate their behavior.
@@ 3 Growing Interest in Data Provenance
Existing norms for tracing AI data provenance have major and increasingly widely acknowledged deficits (Bommasani et al., 2023b; Longpre et al., 2023a). Popular AI systems do not disclose even basic information about their training data (e.g., ChatGPT, Bard, Llama 2). The pace of innovation has prompted community calls for more systematic (Gebru et al., 2021; Bender & Friedman, 2018; Mitchell et al., 2019) and extensive (Sambasivan et al., 2021; Rogers, 2021; Bandy & Vincent, 2021; Dodge et al., 2021; Longpre et al., 2023a) data documentation. However, these calls have resulted in uneven adoption and adherence. Documentation issues remain particularly acute for so-called “datasets-of-datasets”, massive collections of hundreds of datasets where the original provenance information is often neglected or lost due to the lack of standard structures. Meanwhile, practitioners have called for greater data transparency (Bommasani et al., 2023b; Narayanan & Kapoor, 2023), for data supply chain and ecosystem monitoring (Bommasani et al., 2023a; Competition and Markets Authority, 2023; Khan, 2021), for content authenticity verification (Rosenthol, 2022), for detailed provenance tracing on behalf of reproducible, explainable, and trustworthy AI systems (Kale et al., 2023; Werder et al., 2022), and specifically for a standardized database to document trustworthy data (Lohr, 2023; Data Nutrition Team, 2021; Longpre et al., 2023a).
Regulators and lawmakers in many countries have also shown interest in these objectives. The U.S. and EU have taken significant steps toward data transparency and both the EU AI Act and President Biden’s recent executive order on AI [^3] https://www.federalregister.gov/d/2023-24283. include provisions related to transparency, provenance, and the need to thoroughly understand AI models’ inputs (See Section 4.2). The EU Act in particular spells out specific requirements for providers of foundation models related to training data provenance. In addition, several recent U.S. congressional bills [^4] https://www.congress.gov/bill/118th-congress/senate-bill/3312.555https://www.blumenthal.senate.gov/imo/media/doc/09072023bipartisanaiframework.pdf propose regulatory regimes for AI that would require data transparency. A voluntary code of conduct put forward by Canadian authorities (Government of Canada, 2023) calls on model developers to publish training data descriptions while UN bodies have recommended international regulations on data rights that enshrine transparency (United Nations, 2023).
This clear interest from both the researchers and lawmakers motivates our work on unified frameworks for data provenance and transparency. While such standards do not address AI risks directly, they are a key prerequisite to assessing risks and help foster more responsible AI development.
@@ 4.1 Provenance and copyright
A robust data provenance standard could help address many important issues related to the use of copyrighted material for AI training. For both pretraining and finetuning, a standard data provenance framework can help mitigate legal risks and aid in the enforcement of copyright interests. Copyright infringement hinges on access to protected works [^8] Sid & Marty Krofft Television Prods., Inc. v. McDonald’s Corp., 562 F.2d 1157, 1172 (9th Cir. 1977); Art Attacks Ink, LLC v. MGA Entm’t Inc., 581 F.3d 1138, 1143 (9th Cir. 2009) and thus knowing what datasets were used to train a model and what works are contained in these datasets is critical to assessing copyright issues. As discussed, training data is often mixed and repackaged, which complicates the task of precisely identifying what data was used to train a particular model. A robust framework not only helps creators assert their rights when generated outputs infringe on their copyright, but also helps model developers tune their models to avoid this infringement in the first place. Meanwhile, for finetuning and other curated datasets, a data provenance standard can ensure that model developers have access to accurate license information, making it easier to comply with relevant restrictions.
@@ 4.2 AI Regulation
Both the EU AI Act and President Biden’s recent Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence directly and indirectly highlight the need for transparency for AI systems. Both texts require clear communication of the limitations of AI systems to consumers. The AI Act requires the disclosure of relevant information about training, validation, and testing datasets for high-risk AI systems and a summary of copyright protected training data used in foundation models. The technical specifications in the Act include specific data provenance information such as how the data was obtained, labeled, and processed. Meanwhile, the Executive Order encourages regulatory agencies to emphasize transparency requirements for AI models to protect consumers.
@@ 4.3 The role of lawmakers in encouraging responsible AI practices
This paper is a call to action for dataset creators, model developers, researchers, and lawmakers. By understanding the nature of AI ecosystems lawmakers can create incentives that encourage better documentation of new datasets and audits of existing data. While the term “transparency” is often ill-defined in AI regulation, regulators could leverage transparency obligations to encourage model developers to record information about the datasets that have been used to train models. In addition, policymakers could provide funding for research related to data provenance. Today, there are perverse incentives that prevent many companies from disclosing information about their datasets as doing so may increase the probability of legal action. Legal authorities could consider providing safe harbor to organizations that provide necessary information about their datasets to regulators and the public.
@@ 5.3 Dataset Provenance Standards
Datasheets, Statements & Cards. Datasheets (Gebru et al., 2021), data statements (Bender & Friedman, 2018), and data cards (Pushkarna et al., 2022) propose documentation standards for AI datasets. Each of these standardizes documentation of AI dataset creators, annotators, language and content composition, innate biases, collection and curation processes, uses, distribution, and maintenance. Though unevenly adopted, these efforts are widely recognized for improving scientific reproducibility, and responsible AI (Boyd, 2021).
@@ 6 Elements of a Data Provenance Standard
    Verifiable: The metadata can be verified, or its reliability assessed. Although metadata will inevitably contain errors, editing systems (like the one used by Wikipedia), or provenance confirmations created by data creators help to provide transparency and consistency.
@@ 7 Key lessons for data provenance
Existing data provenance solutions are piece-meal and without a robust, well-resourced data provenance framework, developers will struggle to accurately identify and evaluate the safety, copyright implications, and relevance of datasets from a dizzying array of possibilities. Data creators will similarly struggle to identify how and where their content is. Without dataset provenance standards and documentation, creating such a framework will become increasingly difficult and ultimately untenable. While each existing solution provides important insights into the data ecosystem, a robust framework to attach metadata to datasets is needed to track how datasets are mixed, complied and used. We advocate for a set of actions that different stakeholders can take to make data authenticity, consent, and provenance more robust to future challenges.

### v2

@@ Abstract
New capabilities in foundation models are owed in large part to massive, widely-sourced, and under-documented training data collections. Existing practices in data collection have led to challenges in tracing authenticity, verifying consent, preserving privacy, addressing representation and bias, respecting copyright, and overall developing ethical and trustworthy foundation models. In response, regulation is emphasizing the need for training data transparency to understand foundation models’ limitations. Based on a large-scale analysis of the foundation model training data landscape and existing solutions, we identify the missing infrastructure to facilitate responsible foundation model development practices. We examine the current shortcomings of common tools for tracing data authenticity, consent, and documentation, and outline how policymakers, developers, and data creators can facilitate responsible foundation model development by adopting universal data provenance standards.
@@ 1 The Need for Data Provenance
In the last decade, data from across the web, such as news, social media, and encyclopedias, has become a vital resource in general-purpose, generative AI consumer technologies [^1] This work is scoped specifically to AI technologies often called foundation models that are (a) generative, meaning they can create potentially inauthentic text, images or other content, and (b) are typically general-purpose, and therefore tend to scrape vast, diverse distribution of web data, that are under-documented, curated, and checked for consenting use. like GPT-4, Midjourney, and Whisper. These technologies, some with over 100 million weekly users (Malik, 2023), have already begun to catalyze innovation (Brynjolfsson et al., 2018) and scientific inquiry while starting to affect wide swaths of the economy and many everyday consumers (Bommasani et al., 2021). These models are trained on diverse compilations of text, image, and audio data scraped from the web, synthetically generated, or hand-curated. The resulting race to scrape, secure, and mass-produce massive collections of loosely structured data has come with consequences.
Current practices include widely sourcing and bundling data without tracking or vetting their original sources (Longpre et al., 2023a), creator intentions (Chayka, 2023), copyright and licensing status (Bandy & Vincent, 2021), or even basic composition and properties (Dodge et al., 2021). The lack of transparency on this metadata and public infrastructure available to track it has led developers into ethical and legal challenges.
@@
These issues have motivated new data infrastructure and frameworks (which we survey in Section 5) to overcome challenges in sourcing training data responsibly. We isolate several tools in the existing ecosystem for foundation model data management that we identify as commonly used or cited in the AI literature. However, while these tools offer promising solutions, we find they neglect key aspects of the problems, lack interoperability with parallel standards/tools, or haven’t yet succeeded in reaching widespread adoption (Section 5.5 and Section 6).
This position paper argues that no complete system for data provenance exists despite a multitude of solutions to different elements of the problem. A unified framework dedicated to the structured documentation of data properties is needed. This requires action from multiple stakeholders:
@@
To respond to this urgent need, we outline the importance of data provenance for different stakeholders (Section 2), recognition of the problem (Section 3), and legal and regulatory considerations (Section 4). Moreover, we investigate the shortcomings in existing data infrastructure and highlight how the current set of solutions to trace data authenticity, consent, and provenance have limitations and trade-offs (Section 5). Finally, we outline recommendations to facilitate a standardized library for tracing critical data features, emphasize the need for interoperable data documentation standards, and address challenges facing informed and responsible AI (Section 6 and Section 7).
@@ 2.1 Creators: Protecting Rights and Avoiding Harms
The unsettled legal status of AI training data (Congressional Research Service, 2023) has led to some initial compensation proposals (Cheng, 2023), over which creators have little control. Creators are also rarely credited as having contributed training data, despite the occasional artist signature or watermark that slips through into generated content (Vincent, 2023a). This lack of data transparency and legal clarity has led to several lawsuits against leading AI companies (Vincent, 2022; Small, 2023; Brittain, 2023; Vincent, 2023b) and to calls from creators and publishers for strong transparency requirements around AI data (Cooke, 2023; David, 2023b; Bommasani et al., 2023b, 2024).
@@ 2.2 Developers: Informed Data Use
AI developers have an acute interest in data and its provenance for model performance, model behavior, and for anticipating limitations and risks. While model performance generally improves with more data, the quality and diversity of data are also critical factors for reliable performance (Kaplan et al., 2020; Hoffmann et al., 2022). Model behavior tends to emulate the structure and composition of data at both the pretraining and finetuning stages (Chung et al., 2022; Longpre et al., 2023b). For these reasons, developers curate specialized pretraining corpora for scientific writing (Lewkowycz et al., 2022), code (Li et al., 2022), biomedical content (Singhal et al., 2023) and legal works (Henderson et al., 2022). As such, information about data sources and their properties informs AI model training. Data documentation, provenance and analysis tooling have proven particularly important for helping practitioners understand very large datasets (Bender & Friedman, 2018; Gebru et al., 2021). Existing datasets suffer from discoverability of issues related to biased or sexual content (David, 2023a)), private data (Nasr et al., 2023), copyright infringing data (Min et al., 2023), or non-commercially restricted data (Heath, 2023). Lastly, reproducibility and scientific progress more broadly are accelerated by data transparency and structured documentation (Kapoor & Narayanan, 2022).
@@ 2.4 Broader Impacts: Reducing Risk and Fighting Bias
Societal risks from AI run the gamut from privacy violations (Maiberg, 2023) and exposure of personally identifying information, to systemic economic impacts and job displacement (Autor et al., 2022; Gruetzemacher et al., 2020), to bias and discriminatory behavior (Kurita et al., 2019; Kapoor et al., 2024). These risks are fundamentally tied to the data that models are trained on (Longpre et al., 2023b), which, together with the context and affordances of their applications, dictate their behavior.
@@ 3 Growing Interest in Data Provenance
Existing norms for tracing AI data provenance have major and increasingly widely acknowledged deficits (Bommasani et al., 2023b; Longpre et al., 2023a). Popular AI systems do not disclose even basic information about their training data (e.g., ChatGPT, Bard, Llama 2). The pace of innovation has prompted community calls for more systematic (Gebru et al., 2021; Bender & Friedman, 2018; Mitchell et al., 2019) and extensive (Sambasivan et al., 2021; Rogers, 2021; Bandy & Vincent, 2021; Dodge et al., 2021; Longpre et al., 2023a) data documentation. However, these calls have resulted in uneven adoption and adherence. Documentation issues remain particularly acute for so-called “datasets-of-datasets”: massive collections of hundreds of datasets where the original provenance information is often neglected or lost due to the lack of standard structures. Meanwhile, practitioners have called for greater data transparency (Bommasani et al., 2023b; Narayanan & Kapoor, 2023), for data supply chain and ecosystem monitoring (Bommasani et al., 2023a; Competition and Markets Authority, 2023; Khan, 2021), for content authenticity verification (Rosenthol, 2022), for detailed provenance tracing on behalf of reproducible, explainable, and trustworthy AI systems (Kale et al., 2023; Werder et al., 2022), and specifically for a standardized database to document trustworthy data (Lohr, 2023; Data Nutrition Team, 2021; Longpre et al., 2023a).
Regulators and lawmakers in many countries have also shown interest in these objectives. The U.S. and EU have taken significant steps toward data transparency and both the EU AI Act and President Biden’s recent executive order on AI [^3] https://www.federalregister.gov/d/2023-24283. include provisions related to transparency, provenance, and the need to thoroughly understand AI models’ inputs (See Section 4.2). The EU Act, in particular, spells out specific requirements for providers of foundation models related to training data provenance. In addition, several recent U.S. congressional bills [^4] https://www.congress.gov/bill/118th-congress/senate-bill/3312.555https://www.blumenthal.senate.gov/imo/media/doc/09072023bipartisanaiframework.pdf propose regulatory regimes for AI that would require data transparency. A voluntary code of conduct put forward by Canadian authorities (Government of Canada, 2023) calls on model developers to publish training data descriptions while UN bodies have recommended international regulations on data rights that enshrine transparency (United Nations, 2023).
This clear interest from both the researchers and lawmakers motivates our work on unified frameworks for data provenance and transparency. While such standards do not address AI risks directly, they are a key prerequisite to assessing risks and helping foster more responsible AI development.
@@ 4.1 Provenance and copyright
A robust data provenance standard could help address many important issues related to the use of copyrighted material for AI training. For both pretraining and finetuning, a standard data provenance framework can help mitigate legal risks and aid in the enforcement of copyright interests. Copyright infringement hinges on access to protected works [^8] Sid & Marty Krofft Television Prods., Inc. v. McDonald’s Corp., 562 F.2d 1157, 1172 (9th Cir. 1977); Art Attacks Ink, LLC v. MGA Entm’t Inc., 581 F.3d 1138, 1143 (9th Cir. 2009) and thus knowing what datasets were used to train a model and what works are contained in these datasets is critical to assessing copyright issues. As discussed, training data is often mixed and repackaged, which complicates the task of precisely identifying what data was used to train a particular model. A robust framework not only helps creators assert their rights when generated outputs infringe on their copyright, but also helps model developers tune their models to avoid this infringement in the first place. Meanwhile, for finetuning and other curated datasets, a data provenance standard can ensure that model developers have access to accurate license information, making it easier to comply with relevant restrictions. We note that this analysis focuses on the U.S. fair use principle and that other jurisdictions have different approaches to copyright (e.g. the EU’s Text and Data Mining Exception and Singapore’s Copyright Act 2021).
@@ 4.2 AI Regulation
Both the EU AI Act and President Biden’s recent Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence directly and indirectly highlight the need for transparency for AI systems. Both texts require clear communication of the limitations of AI systems to consumers. The AI Act requires the disclosure of relevant information about training, validation, and testing datasets for high-risk AI systems and a summary of copyright protected training data used in foundation models. The technical specifications in the Act include specific data provenance information such as how the data was obtained, labeled, and processed. Meanwhile, the Executive Order encourages regulatory agencies to emphasize transparency requirements for AI models to protect consumers. We note that privacy regulations like the EU’s GDPR are generally less relevant in the context of finetuning data which does not generally fall under the definition of personal data per Art. 4(1) of GDPR.
@@ 4.3 The role of lawmakers in encouraging responsible AI practices
This paper is a call to action for dataset creators, model developers, researchers, and lawmakers. By understanding the nature of AI ecosystems, lawmakers can create incentives that encourage better documentation of new datasets and audits of existing data. While the term “transparency” is often ill-defined in AI regulation, regulators could leverage transparency obligations to encourage model developers to record information about the datasets that have been used to train models. In addition, policymakers could provide funding for research related to data provenance. Today, there are perverse incentives that prevent many companies from disclosing information about their datasets as doing so may increase the probability of legal action. Legal authorities could consider providing safe harbor to organizations that provide necessary information about their datasets to regulators and the public.
@@ 5.3 Dataset Provenance Standards
Datasheets, Statements & Cards. Datasheets (Gebru et al., 2021), data statements (Bender & Friedman, 2018), and data cards (Pushkarna et al., 2022) propose documentation standards for AI datasets. Each of these standardizes documentation of AI dataset creators, annotators, language and content composition, innate biases, collection and curation processes, uses, distribution, and maintenance. Though unevenly adopted, these efforts are widely recognized for improving scientific reproducibility and responsible AI development (Boyd, 2021).
@@ 6 Elements of a Data Provenance Standard
    Verifiable: The metadata can be verified, or its reliability assessed. Although metadata will inevitably contain errors, editing systems (like the one used by Wikipedia) or provenance confirmations created by data creators help to provide transparency and consistency.
@@ 7 Key Lessons for Data Provenance
Existing data provenance solutions are piece-meal. Without a robust, well-resourced data provenance framework, developers will struggle to accurately identify and evaluate the safety, copyright implications, and relevance of datasets from a dizzying array of possibilities. Data creators will similarly struggle to identify how and where their content is. Without dataset provenance standards and documentation, creating such a framework will become increasingly difficult and ultimately untenable. While each existing solution provides important insights into the data ecosystem, a robust framework to attach metadata to datasets is needed to track how datasets are mixed, complied and used. We advocate for a set of actions that different stakeholders can take to make data authenticity, consent, and provenance more robust to future challenges.